Magnetic Corporation expects dividends to grow at a rate of 19.70% for the next two years. After two years dividends are expected to grow at a constant rate of 06.70% indefinitely. Magnetic’s required rate of return is 08.49% and they paid a $1.91 dividend today. Find the value of Magnetic Corporation’s common stock per share by computing:
       a) Dividend at the end of Year 1:
       b) Dividend at the end of Year 2:
       c) Dividend at the end of Year 3:
       d) Price of stock at end of year 2:
       e) Price of stock today:
Answer #1


Solution: D1=DO*(1+g) = 1.91*1.197 2.28627 2.73667 b) D2=D1*(1+g) = 2.28627*1.197 c) D3=D2*(1+g) = 2.73667*1.067 2.92002 163.